# spacing=nonfrench

### v3

@@ 6. Evaluation
We first present microbenchmark on beam pattern, coverage improvement, and dual-link capabilities of Wall-Street’s (Section˜6.2). We then conduct field studies to evaluate the end-to-end performance in mobile scenarios, comparing Wall-Street against standard 5G SA handover protocol (Section˜6.3).
@@ 6.1. Methodology
We evaluate Wall-Street in both controlled indoor (single-UE) and real-world outdoor (multi-UE) driving scenarios. For indoor settings, there are four gNB nodes and one UE node in a $20\times 30$ m lab, as shown in Fig.˜11(a). To test the UE and Wall-Street under mobility, we place them on a cart moving in a constant speed of approximately 1 km/h. For outdoor settings, there are three gNB nodes on the first floor of a lab facing a parking lot. We mount the surface on the rear window of a SUV vehicle and deploy two UE nodes. UE 1 is on the rear seat (opposite side from surface), and UE 2 is on the cargo area, as shown in Fig.˜10. The vehicle drives at speeds ranging from 5 to 15 km/h. Both the surface and UEs are battery-powered, with UE positions remaining the same for the baseline. Since the gNB nodes use stationary PAAMs that cannot be relocated, our outdoor driving trajectory is limited to 30 m, as depicted in Fig.˜11.
We implement the 5G SA HO protocol as our baseline and use this as a performance benchmark. For both Wall-Street and baseline, the hysteresis parameter $H$ of A3 event is set to $10$ dB with TTT of $150$ ms and the measurement report periodicity of $160$ ms. For the baseline, a total duration of every measurement report takes less than $20$ ms with three neighbor gNBs. Also, we set the maximum in-vehicle distance to $1$ m for $L_{min}$. Our results combine (1) PHY measurements, including RSRP and packet error rate, from COSMOS testbed with real vehicle trajectory, and (2) TCP evaluation using ns-3 simulator driven by PHY traces. In the ns-3 simulator, we increased all outdoor RSRP traces by 10 dB for both the baseline and Wall-Street scenarios.
@@ 6.2. Microbenchmarks
We evaluate Wall-Street’s far-field dual-beam pattern measured with the spectrum analyzer, in-vehicle signal coverage, and dual-beam operations at the physical layer.
Far-field beam patterns. Figure˜12 demonstrates Wall-Street’s beam steering capabilities in four configurations: dual-transmissive mode where both beams pass through the surface (Fig.˜12 (a-b)), and transflective mode where one beam transmits through while another reflects (Fig.˜12 (c-d)). Each configuration shows both uplink and downlink operations, validating bidirectional functionality.
We evaluate two beam angle pairs (-40°/40°, -45°/68°), representing typical handover scenario (see Appendix˜B for angle selections). To measure beam patterns, we place Tx and Rx horn antennas two meters from Wall-Street and sweep the Rx from -70°to 70°. Two transmitters operate on separate frequencies, 26.0 GHz (blue $\triangle$) and 26.1 GHz (red $\triangle$), enabling simultaneous beam measurement. Gray lines mark the target steering angles. Without an anechoic chamber, a large portion of signal power diffracts around the surface. We quantify this baseline leakage by disabling Wall-Street (0V, shown as $\blacksquare$) and normalize all measurements to the spectrum analyzer’s noise floor (0 dB reference).
Figure 13. Power allocation between beams at -40°/55°, normalized to baseline. Balanced (50/50) vs. asymmetric (75/25) distribution. Gray lines indicate target angles.
Results show 20-25 dB gain over baseline at target angles, except when a steered beam coincides with the natural specular reflection angle (visible at -40°/40° in Fig.˜12(c) and Fig.˜12(d)). Gain decreases at wider steering angles. We use identical voltage configurations for both uplink and downlink (e.g., Fig.˜12(a) and Fig.˜12(b)), confirming angular reciprocity that allows fast UL/DL switching during handover.
Beam power ratio. Figure˜13 shows dynamic power allocation between beams at -40°/55°. All measurements are normalized against the baseline (surface off) for visibility. Shifting from balanced (50/50) to asymmetric (75/25) distribution increases the first beam (blue) while decreasing the second beam (red) at target angles. Both beams maintain target angles (gray lines) and only relative power changes.
In-vehicle mmWave coverage. We evaluate Wall-Street’s ability to overcome vehicle body blockage by mounting it on a vehicle’s rear window with a gNB positioned 30 m away. For each in-vehicle location, we measure maximum achievable RSRP by exhaustively searching all beam combinations. In Figure˜14, nearly 50% of the vehicle interior experiences complete signal outage without Wall-Street. On the other hand, Wall-Street ensures outage-free coverage inside the vehicle and provides signal strength improvement of over 12 dB.
@@
Figure 15. Signal strength and beam angles during vehicle movement. Upper: RSRP; middle: corresponding gNB beam angles; lower: UE/surface angles.
Beam tracking under mobility. Figure˜15 shows RSRPs and corresponding beam angles for gNB and UE/surface as the vehicle moves along a 30-meter outdoor path for UE 1 (Fig.˜15(a)) and UE 2 (Fig.˜15(b)). Wall-Street provides significantly higher signal strength overall. This improvement becomes particularly evident for UE 2. Without Wall-Street, UE 2 loses signals 70-80% of the time despite three available gNBs, whereas with Wall-Street, it receives consistent power from either gNB 1 or gNB 3 throughout the entire path.
@@
Multi-link operations. Figure˜16 evaluates Wall-Street’s ability to handle multiple simultaneous beams at various indoor locations. Figure˜16(a) shows baseline single-link performance without Wall-Street. The upper heatmaps demonstrate best-case RSRP when UE beams are optimally aligned with each gNB. The lower heatmaps show worst-case RSRP when beam is misaligned. Figure˜16(b) demonstrates Wall-Street’s dual-link capabilities. The upper heatmaps show dual-link signal reception where Wall-Street splits power equally between gNB1 and gNB4. While individual RSRPs are about 3dB lower than optimal alignment due to power division, both links remain active. The lower heatmaps validate our RSRP inference method (Section˜4.2.3) by showing signal strengths calculated from reflected measurements.
@@
Figure 17. Throughput and round trip time (RTT) achieved by four congestion control algorithms. The right and lower edge of the box represents the 10% percentile, respectively. The left and upper edge give the 90% percentiles. The intersection point of the horizontal and vertical error bar represents the median value.
6.3. End-to-end Throughput and Delay
We evaluate end-to-end performance of Wall-Street delivering bulk TCP data flows. To quantify performance, we calculate throughput as the total data bits divided by the duration, measured every 100 ms. We average the round-trip-time (RTT) of packets over every 100 ms.
@@ 6.3. End-to-end Throughput and Delay
TCP performance under various settings. We evaluate the throughput and RTT of Wall-Street and the baseline under indoor and outdoor settings using four congestion control algorithms: CUBIC, Vegas, HighSpeed, and NewReno. In the indoor setting (Fig.˜17(a)), Wall-Street enhances average throughput by at least 20% and reduces delay by 20-80% compared to the baseline. For the outdoor multi-user scenario (Fig.˜17(b)), Wall-Street yields more significant improvements. Throughput improves by up to 78% (68% on average for UE 1 and 60% on average for UE 2) and RTT reduction by up to 34% (14% on average for UE 1 and 5% on average for UE 2). Notably, in the baseline setup, UE 2 (located in the cargo area) experiences lower throughput and higher RTT than UE 1 due to car body blockage.
@@
TCP performance over time. Figure˜18 divides the 20-second outdoor trajectory into 10 intervals, showing median throughput and delay for UE 1 (upper) and UE 2 (lower). Vertical dashed lines mark handover events. Wall-Street demonstrates three key advantages over baseline. First, Wall-Street maintains higher throughput and lower latency throughout the trajectory, with UE 1 achieving higher throughput improvement and UE 2 maintaining connectivity where baseline fails. Second, at the 6s mark, Wall-Street executes a single batch handover for both UEs simultaneously. This contrasts with baseline, where UE 1 and UE 2 trigger separate handovers at different times, causing individual service interruptions and increased signaling overhead. Lastly, baseline suffers from ping-pong handovers between gNB 1 and gNB 3 and a complete connection loss for UE 2 at 12s following a failed handover. Wall-Street eliminates both issues through stable decision-making and make-before-break execution.
Figure 18. Throughput and delay of UE 1 (upper) and UE 2 (lower) achieved without Wall-Street (left) and with Wall-Street (right) as the vehicle is moving along the same trajectory. Vertical dashed lines indicate the handover event.
Figure 19. Impact of speed from 5 km/h to 15 km/h: throughput, RTT, HO count, and total outage duration (left to right).
@@
Impact of vehicle speed. Figure˜19 compares Wall-Street’s performance against the baseline at driving speeds from 5 km/h to 15 km/h. Wall-Street consistently outperforms the baseline across most metrics and speed scenarios. It improves user throughput by 70% at 5 km/h, 27% at 10 km/h, and 8.4% at 15 km/h. It also makes the network connection more reliable by reducing the number of handovers by an average of 37% and delay by an average of 11% and cutting connection outages by an average of 64%. Wall-Street maintains the same low number of handovers regardless of speed because it makes one conservative decision for all users simultaneously based on multiple measurements, preventing the unnecessary handovers that happens the baseline. However, performance gains diminish at higher speeds due to a limitation of our current implementation where the system cannot complete handover preparation quickly enough. We discuss potential solutions to accelerate the scanning in Section˜7.
@@
Impact of handover preparation. Figure˜20 illustrates sequence number progression during the outdoor experiment, with the left subfigure showing the entire duration and the right subfigure focusing on the handover preparation phase. To enhance visibility of trends, sequence numbers are subjected to a modulo operation. Wall-Street allows a steeper increase in sequence numbers compared to the baseline, indicating that more packets are successfully delivered. During handover preparation, the baseline’s sequence numbers show irregular progression and a lower slope due to service interruptions as the UE stops communicating with the source gNB to detect SSB bursts from neighboring gNBs. While SSB detection takes under 20 ms, its effects persist for 50 ms, significantly disrupting baseline packet delivery. In contrast, Wall-Street maintains steady sequence number progression throughout the handover preparation phase, demonstrating more stable communication.
@@
Impact of make-before-break handover. Figure˜21 evaluates make-before-break performance using real-time packet error rate (PER) measurements during a 25-second drive at 5 km/h. Both gNB1 and gNB3 transmit duplicate packets (identified by sequence numbers), enabling per-link and combined PER calculation over time.
The baseline setup faces a fundamental limitation where UE 1 can only align with one gNB despite both transmitting. We configure UE 1’s beam toward gNB 3 while both gNBs beamform toward UE 1. This single-alignment constraint means gNB 1’s packets rarely reach the UE, negating make-before-break benefits.
Wall-Street eliminates this limitation by simultaneously steering both gNB signals to the UE. The upper plots show individual link PER with gNB1 (red) and gNB3 (blue), while lower plots show combined PER when receiving packets from both source (black).
Without Wall-Street, gNB 3 provides relatively good PER but gNB 1 remains mostly unusable, rarely benefiting from make-before-break mechanism. With Wall-Street, both links operate at moderate PER due to power splitting, but their combination achieves low PER throughout the trajectory. This validates that Wall-Street transforms make-before-break from a theoretical capability into practical reliability.
Figure 21. Make-before-break impact: real-time packet error rates (PERs) of a link from each gNB (upper) and that of two links with duplicated packets from both gNBs (lower).
7. Discussion
In this section, we discuss Wall-Street’s current limitations and potential solutions.
@@ 7. Discussion
Control Channel. Our prototype uses an out-of-band sub-6 GHz link to integrate with the COSMOS roadside testbed, as the testbed requires all components to be connected through the same local network. However, Wall-Street only requires a few control bits to index its pre-computed voltage codebook. In deployment, this control information can move in-band. The gNB can embed low-rate control data directly into the mmWave Synchronization Signal Blocks (SSBs) it already transmits, using simple modulation schemes such as Frequency-Shift Keying (FSK) on a subcarrier. The surface would then need only a low-power detector synchronized to the SSB periodicity to decode these commands, eliminating the need for an active modem entirely.
@@
Reducing Alignment Overhead. RIS-assisted roadside networking faces significant beam alignment overhead across the gNB, surface, and UE. We can reduce this complexity through two approaches. First, vehicles’ existing GPS systems provide rough location estimates that allow the gNB to narrow its transmit beam search space, effectively eliminating one search dimension during the beam alignment. We note that we may use GPS only for accelerating beam alignment, but not for handover decisions as location alone cannot capture environmental blockages or vehicle body effects that determine actual signal quality between the base station and UEs. Second, the surface can operate as a reflective probe during the initial alignment. It can reflect and steer the gNB’s SSB transmissions while the gNB listens for the strongest return, taking $O(N^{2})$ to determine their mutual orientation. The surface can then reflect and steer UE transmission toward the UE with another $O(N^{2})$ search. This two-stage approach could reduce the complexity of initial alignment.

### v4

@@ 6. Evaluation
We present microbenchmarks on beam patterns, coverage improvement, and dual-link capabilities of Wall-Street (Section˜6.2). We then conduct field studies to collect traces from the COSMOS testbed under driving scenarios. Finally, we input these PHY traces into the ns-3 simulator to analyze TCP throughput and latency performance (Section˜6.3).
@@ 6.1. Methodology
We evaluate Wall-Street in both controlled indoor (single-UE) and real-world outdoor (multi-UE) driving settings. For indoor settings, there are four gNB nodes and one UE node in a $20\times 30$ m lab, as shown in Fig.˜9(a). To test the UE and Wall-Street under mobility, we place them on a cart moving in a constant speed of approximately 1 km/h. For outdoor settings, there are three gNB nodes on the first floor of a lab facing a parking lot. We mount the surface on the rear window of a SUV vehicle and deploy two UE nodes. UE 1 is on the rear seat (opposite side from surface), and UE 2 is on the cargo area, as shown in Fig.˜8. The vehicle drives at speeds ranging from 5 to 15 km/h. Both the surface and UEs are battery-powered, with UE positions remaining the same with and without Wall-Street. Since the gNB nodes use stationary PAAMs that cannot be relocated, our outdoor driving trajectory is limited to 30 m, as depicted in Fig.˜9.
Figure 10. Beam patterns for four steering modes. Red/blue $\triangle$ show measurements at 26.0/26.1 GHz. $\blacksquare$ show baseline (surface turned off). Gray lines indicate groundtruth angles. Left: -40°/40° steering; right: -45°/68° steering.
We implement the standard SA protocol as our baseline. For both Wall-Street and baseline, the hysteresis parameter $H$ of A3 event is set to $10$ dB with TTT of $150$ ms and the measurement report periodicity of $160$ ms. For the baseline, a total duration of every measurement report takes less than $20$ ms with three neighbor gNBs. Also, we set the maximum in-vehicle distance to $1$ m for $L_{min}$. Our results combine (1) PHY measurements, including RSRP and packet error rate, from the COSMOS testbed with real vehicle trajectory, and (2) TCP evaluation using ns-3 simulator driven by PHY traces. As RSRP values from the COSMOS testbed serve as relative signal strength indicators defined by the testbed and do not directly map to absolute dBm scales, we applied a 10 dB calibration offset to all traces, aligning them with the absolute dBm scale required by the ns-3 simulator.
@@ 6.2. Microbenchmarks
We evaluate Wall-Street’s far-field dual-beam patterns measured with a spectrum analyzer, in-vehicle signal coverage, and dual-beam performance at the physical layer.
Far-field beam patterns. Figure˜10 demonstrates Wall-Street’s beam steering capabilities in four configurations: dual-transmissive mode where both beams pass through the surface (Fig.˜10 (a-b)), and transflective mode where one beam transmits through while another reflects (Fig.˜10 (c-d)). Each configuration shows both uplink and downlink operations, validating its bidirectional functionality.
We evaluate two beam angle pairs (-40°/40°, -45°/68°). To measure beam patterns, we place Tx and Rx horn antennas two meters from Wall-Street and sweep the Rx from -70°to 70°. Two transmitters operate on separate frequencies, 26.0 GHz (blue $\triangle$) and 26.1 GHz (red $\triangle$), for simultaneous beam measurement. Gray lines indicate the groundtruth steering angles. In the absence of an anechoic chamber, a significant portion of signal power diffracts around the surface. We quantify this background leakage by setting Wall-Street to a uniform 0V state—effectively acting as a specular mirror ($\blacksquare$). We normalize all measurements to the spectrum analyzer’s noise floor (0 dB reference).
Figure 11. Power allocation between beams at -40°/55°, normalized to baseline. Balanced (50/50) vs. asymmetric (75/25) ratio. Gray lines indicate groundtruth angles.
We observe a 20–25 dB gain at ground-truth steering angles. This gain persists except when the steered beam coincides with the natural specular reflection angle, as seen at -40°/40° in Fig.˜10(c) and Fig.˜10(d), where the baseline ($\blacksquare$) itself is high. This gain naturally decreases at wider steering angles. We use identical voltage configurations for both uplink and downlink (e.g., Fig.˜10(a) and Fig.˜10(b)), confirming angular reciprocity that allows fast UL/DL switching during handovers.
Beam power ratio. Figure˜11 shows dynamic power allocation between two beams at -40°/55°. All measurements are normalized against the baseline (surface off) for visibility. Shifting from balanced (50/50) to asymmetric (75/25) ratio increases the first beam (blue) while decreasing the second beam (red) at groundtruth angles. Both beams maintain groundtruth angles and only relative power changes.
@@
In-vehicle mmWave coverage. To evaluate Wall-Street’s ability to mitigate vehicle body blockage, we mounted the surface on the rear window with a gNB located 30 m away. For each interior location, we determined the maximum achievable RSRP through an exhaustive beam search. As shown in Figure˜12, nearly 50% of the vehicle interior experiences complete signal outage without Wall-Street. In contrast, Wall-Street eliminates these outages entirely and provides a signal strength improvement of over 12 dB.
@@
Figure 13. Signal strength and beam angles during outdoor driving experiments (upper: RSRP; middle: corresponding gNB beam angles; lower: UE or surface angles).
Beam tracking under mobility. Figure˜13 shows RSRPs and corresponding beam angles for gNB and UE/surface as the vehicle moves along a 30-meter outdoor path for UE 1 (Fig.˜13(a)) and UE 2 (Fig.˜13(b)). Overall, Wall-Street provides significantly higher signal strength. This becomes particularly evident for UE 2. Without Wall-Street, UE 2 loses signals 70-80% of the time, despite the presence of three gNBs. In contrast, Wall-Street ensures continuous connectivity from either gNB 1 or gNB 3 throughout the entire path.
@@
Multi-link operations. Figure˜14 evaluates Wall-Street’s ability to simultaneously control two beams in an indoor environment. In Fig.˜14(a), we show baseline performance without the surface. The upper heatmaps display the best-case RSRP from gNB 1 (left) or gNB 4 (right) with optimal UE beam alignment, while the lower heatmaps depict the worst-case scenario where beams are misaligned. Figure˜14(b) demonstrates Wall-Street’s dual-link capability. The upper heatmaps reveal simultaneous signal reception where Wall-Street splits power equally between gNB1 and gNB4. Although individual RSRPs are approximately 3 dB lower than single-beam alignment due to this power division, both links remain active. The lower heatmaps demonstrate our RSRP inference method by showing signal strengths calculated from reflected measurements.
6.3. Throughput and Latency
We evaluate throughput and latency performance of Wall-Street delivering bulk TCP data flows. To quantify performance, we calculate throughput as the total data bits divided by the duration, measured every 100 ms. We average the round-trip-time (RTT) of packets over every 100 ms.
@@ 6.3. Throughput and Latency
Figure 15. Throughput and round trip time (RTT) with four congestion control algorithms. The right and lower edge of the box represents the 10% percentile, respectively. The left and upper edge give the 90% percentiles. The intersection point represents the median value.
@@
TCP performance. We evaluate the throughput and Round-Trip Time (RTT) of Wall-Street against the baseline in both indoor and outdoor settings, using four congestion control algorithms: CUBIC, Vegas, HighSpeed, and NewReno. For indoor (Fig.˜15(a)), Wall-Street enhances throughput by at least 20% and reduces delay by 20-80%. Under outdoor multi-user settings (Fig.˜15(b)), Wall-Street yields more significant improvements. Throughput improves by up to 78% (68% on average for UE 1 and 60% on average for UE 2) and delay reduces by up to 34% (14% on average for UE 1 and 5% on average for UE 2). Notably, in the baseline setup, UE 2 (located in the cargo area) suffers from lower throughput and higher RTT due to vehicle body blockage
@@
TCP performance over time. Figure˜16 divides the 20-second outdoor trajectory into 10 intervals, showing median throughput and delay for UE 1 (upper) and UE 2 (lower) over time. Vertical dashed lines denote HO triggers. Wall-Street demonstrates three key advantages over baseline. First, Wall-Street maintains higher throughput and lower latency throughout the trajectory. UE 1 achieves higher throughput improvement, and UE 2 maintains continuous connectivity at the 12s mark where the baseline forces a connection termination. Second, at the 6s mark, Wall-Street executes a collective HO for both UEs. In contrast, the baseline triggers separate HOs, causing individual service interruptions and increased signaling overhead. Lastly, the baseline suffers from ping-pong HOs between gNB 1 and gNB 3. Wall-Street eliminates this instability through robust decision logic.
@@
Impact of vehicle speed. Figure˜17 compares Wall-Street’s performance against the baseline at driving speeds from 5 to 15 km/h. Wall-Street consistently outperforms the baseline across most metrics. Specifically, it boosts user throughput by 70% at 5 km/h, 27% at 10 km/h, and 8.4% at 15 km/h. Reliability also improves significantly as Wall-Street reduces HOs by an average of 37%, lowers latency by 11%, and cuts connection outages by 64%. Wall-Street maintains a consistently low HO count regardless of speed because it makes one conservative decision for all UEs based on multiple measurements, preventing the unnecessary HOs observed in the baseline. We note that performance gains diminish at higher speeds due to latency constraints in our current implementation, which limit the speed of HO preparation.
Figure 16. Throughput and RTT over time without Wall-Street (left) and with Wall-Street (right) for UE 1 (upper) and UE 2 (lower) as the vehicle moves along the same trajectory. Vertical dashed lines indicate HO triggers.
Figure 17. Impact of speed from 5 to 15 km/h: throughput, RTT, HO count, and total outage duration (left to right).
@@
Impact of handover preparation. Figure˜18 tracks the progression of sequence numbers during the outdoor experiment. The left subfigure covers the full duration, while the right focuses specifically on the HO preparation phase. We apply a modulo operation to sequence numbers to visualize trends. Wall-Street maintains a continuous, steep slope, demonstrating successful packet delivery during HO preparation. The baseline, however, shows irregular progression and a lower slope because the UE stops data transmission to scan for neighboring gNB scanning. While SSB detection takes under 20 ms, its effects persist for 50 ms, significantly disrupting baseline packet delivery.
@@
Impact of make-before-break handover. Figure˜19 evaluates Make-Before-Break (MBB) performance via real-time PER measurements over a 25-second drive at 5 km/h. In this setup, gNB 1 and gNB 3 transmit duplicate packet streams for both the Wall-Street and baseline scenarios. We extract both per-link and combined PER metrics. The baseline setup is inherently limited by the UE’s single-beam constraint. Even though both gNBs are transmitting, the UE can only align its beam with one (here, gNB 3). Consequently, packets from gNB 1 rarely reach the UE, effectively negating the benefits of MBB. In contrast, Wall-Street simultaneously steer both gNB signals to the UE. While individual links (red/blue) operate at moderate PER due to power splitting, their combined reception (black) achieves consistently low PER.
Figure 19. Real-time make-before-break performance under mobility: PER of individual links from each gNB (upper), and combined PER using duplicate packets from two gNBs (lower). Data reflects real-time driving data.
[…]
Figure 20. UE power measurement: power increases from 66 W to 95 W during the search process.
7. Limitations and Discussions
To integrate our system with cellular infrastructure, we make several non-standard compliant assumptions. In this section, we outline each assumption and discuss potential solutions.
gNB-side Measurement. Our system assumes the serving gNB can measure reflected signals from a neighbor gNB while serving the UE. To adopt our system under TDD deployment, we can synchronize gNBs using inter-gNB coordination. We can use the standard Xn interface, which allows gNBs to exchange configuration data, including TDD slot formats and SSB periodicity. The serving gNB can acquire the neighbor gNB’s broadcast schedule relative to its own internal clock and align an uplink slot to overlap with the neighbor’s SSB transmission. During this slot, the serving gNB is in receive mode and measures the Wall-Street reflection. We can also implement our system into an O-RAN architecture, offloading this coordination logic to the O-RAN Near-Real-Time RAN Intelligent Controller (Near-RT RIC).
@@ 7. Limitations and Discussions
Control Channel. Our system uses an out-of-band sub-6 GHz link to integrate with the COSMOS roadside testbed, as the testbed requires all components to be connected through the same local network. Here, Wall-Street receives a few control bits to index its pre-computed voltage codebook. In deployment, this control information can move in-band. The gNB can embed low-rate control data directly into the mmWave SSB it already transmits, using simple modulation schemes such as Frequency-Shift Keying (FSK) on a subcarrier. The surface would then need only a low-power detector synchronized to the SSB periodicity to decode these commands, eliminating the need for a sub-6 modem entirely.
@@
Reducing Alignment Overhead. Initial beam alignment requires high overhead across the gNB, surface, and UE. We can reduce this complexity in two ways. First, we leverage vehicle GPS for coarse location estimation, allowing the gNB to narrow its transmit beam search sector. We note that we may use GPS for accelerating beam search, but not for handover decisions as it cannot capture environmental blockages or vehicle body effects that determine actual signal quality. Second, the surface can operate as a reflective probe. It can reflect and steer the gNB’s signals while the gNB listens for the strongest return, taking $O(N^{2})$ to determine their mutual orientation. The surface can then reflect and steer UE signals toward the UE with another $O(N^{2})$ search.
Deployment Feasibility. We propose mounting the surface on the vehicle’s interior sunroof or glass ceiling. This placement aligns with the arrival angle of downlink signals while protecting the hardware from rain and snow. Our link-budget analysis (Appendix B) confirms that a compact $10\times 20$ cm surface provides sufficient signal strength for both user connectivity and cell search under real-world settings.
Scalability. We address scalability across two dimensions:
  • $\bullet$
    Interference: As the number of vehicles increases, surface reflections could interfere with nearby networks or outdoor users. Since the metamaterial surface use tight element spacing, precise phase control could further nullify interference. Also, because gNBs transmit downwards onto the roof-mounted surface, reflections are directed upward, minimizing leakage to ground-level outdoor users.
  • $\bullet$
    UE contention: Wall-Street uses time-division to serve multiple users, which causes contention as the number of UEs increases. We may partition the surface into multiple sub-areas to serve multiple users simultaneously when not performing handovers. A MAC-layer scheduler can dynamically manage these resources, allocating larger surface areas to users with higher throughput demands.
Complex Mobility. Our outdoor experiments were constrained by the testbed’s fixed topology and control network access. Real-world driving involves unpredictable turns and varying speeds. At a distance of 150 meters with a standard 15° beamwidth, the mmWave beam footprint on the ground is approximately 40 m wide. This is sufficient to cover a bus even during quick turns. We may also use GPS data for beam tracking to add robustness against rapid speed changes.